I need some help with some algebra word problems. I can't quite figure out the answers.
If you could help me i would greatly appreciate it.
The difference between the width and the length of a field is 12 m. Find its dimensions if a fence 160 m in length encloses it.
The length of a rectangle is found by doubling thw width and subtracting 3 m. Find its dimensions if the perimeter is 84 m.

Answer by ankor@dixie-net.com(12701)

"The difference between the width and the length of a field is 12 m."
This means: If width = w; then (w+12) = length
:
Find its dimensions if a fence 160 m in length encloses it.
:
Perimeter:
2L + 2W = 160
:
Substitute (w+12) for L, solve for w
2(w+12) + 2w = 160
2w + 24 + 2w = 160
4w = 160 - 24
4w = 136
w = 136/4
w = 34 m is the width
:
34 + 12 = 46 m is the length:
:
Check using the perimeter:
2(46) + 2(34) =
92 + 68 = 160
:
:
The length of a rectangle is found by doubling thw width and subtracting 3 m.
This means L = 2W - 3
:
Find its dimensions if the perimeter is 84 m.
Substitute (2w-3) for L in the perimeter equation and solve for w
:
Do it exactly like the 1st one.
You can email me if you have difficulty with this